# Scalable Distributed Nonlinear Control Under Flatness-Preserving Coupling

### v1

@@ I-A Related Work
Differential flatness has long been exploited in the control of single-agent systems, as well as in consensus and formation control for multi-agent systems whose subsystems are completely decoupled (Mondal and Tsourdos, 2022, 2025; Van Parys and Pipeleers, 2017). However, for dynamically coupled flat subsystems, flatness of the joint system is not guaranteed, nor is it clear how to construct the joint diffeomorphism in general. Menara et al. (2020) provide a sufficient condition for the flatness of complex networks, but require solving a partial differential equation for the flat outputs. Bidram et al. (2013, 2014) apply the closely-related notion of feedback linearization to a network of dynamically coupled power generators, but assume that the coupling is quasi-static and thus can be approximated as a constant in the diffeomorphism. In contrast, focusing on pure-feedback systems, we identify a class of dynamic couplings where the joint system is provably flat and explicitly construct its flat outputs and diffeomorphism, without assuming quasi-static coupling.
@@ III Flatness-Preserving Coupling
We denote the coupling term affecting the $i^{\textrm{th}}$ subsystem as $\Delta^{i}$, and further, its effect on the $j^{\textrm{th}}$ substate as $\Delta^{i}_{j}$, such that
@@
A coupling term ${\Delta}(\mathbf{x})$ is lower-triangular on $\mathcal{X}$ if its effect on the $j^{\textrm{th}}$ order substate of the $i^{\textrm{th}}$ subsystem $\Delta^{i}_{j}$ is smooth and can be expressed as
@@
While lower-order states from any subsystem may affect the $i$-th subsystem’s states of the same or higher order, the dependence on $\mathbf{x}^{<i}_{j+1}$ in (6) implicitly imposes an acyclic structure: higher-indexed subsystems may be affected by $(j+1)^{\textrm{th}}$ order states from lower-indexed ones, but not vice versa. For simplicity, we present Definition 1 and the subsequent results for the canonical case where this acyclic ordering is fixed across $\mathcal{X}$ and coincides with the subsystem indices. We emphasize that the acyclicity restriction only applies to $(j+1)^{\textrm{th}}$ order dependencies. Further, in Section IV, we analyze a special case where such dependencies are absent and thus no acyclicity is required. We leave extending our results to a case where the ordering is state-varying to future work. In Figure 1, we show an illustration of these structural assumptions for a specific choice of $i$ and $j$.
@@
for all $(\mathbf{x},\mathbf{u})\in\mathcal{X}$} is known [^5] The existence of such maps are guaranteed by Assumption 1 and the implicit function theorem. (See Yang et al. (2025) for a proof.).
@@ V Distributed Tracking Controller
where $K^{i}\in\mathbb{R}^{m_{i}\times rm_{i}}$ is a stabilizing feedback gain, which can be chosen, e.g., via pole placement. We can then compute the control input $\mathbf{u}^{i}$ by evaluating the flatness diffeomorphism $\Psi^{i}$, which depends on flat state $\mathbf{z}^{i}$, virtual input $\mathbf{v}^{i}$, and $(\mathbf{z}^{j},\mathbf{v}^{j})$ of an appropriate subset of the other subsystems. By the results in Section IV, subsystem $i$ only needs to gather $(\mathbf{z}^{j},\mathbf{v}^{j})$ from subsystems $j\in\mathbf{Anc}_{\mathcal{G}(\mathbf{x})}(i)(\mathbf{x})$ to evaluate $\Psi^{i}$. Further, when the coupling is strongly lower-triangular, this dependency is further localized to its $r$-hop neighbors, $\mathcal{N}^{r}_{\mathcal{G}(\mathbf{x})}(i)$. This allows for efficient distributed implementation as the information sharing between the systems does not scale with the number of subsystems $N$. The overall pipeline is summarized in Algorithm 1.
@@ VII Conclusion
We proposed a distributed controller for networks of flat, pure-feedback systems under dynamic coupling. We identified a class of lower-triangular couplings that preserve the flatness for the joint system and allow the joint flatness diffeomorphism to be explicitly constructed. We analyzed and exploited the sparsity pattern of the flatness diffeomorphism to design distributed controllers under communication constraints. We validated the framework in simulation for planar quadrotors coupled by downwash drag, showing that our flatness-based distributed controller achieves accurate tracking. Future work includes hardware experiments as well as relaxing the lower-triangular constraints on flatness-preserving couplings.

### v2

@@ I-A Related Work
Differential flatness has long been exploited in the control of single-agent systems, as well as in consensus and formation control for multi-agent systems whose subsystems are completely decoupled (Mondal and Tsourdos, 2022, 2025; Van Parys and Pipeleers, 2017). However, for dynamically coupled flat subsystems, flatness of the joint system is not guaranteed, nor is it clear how to construct the joint diffeomorphism in general (Nicolau and Iovine, 2022). Menara et al. (2020) provide a sufficient condition for the flatness of complex networks, but require solving a partial differential equation for the flat outputs. Bidram et al. (2013, 2014) apply the closely-related notion of feedback linearization to a network of dynamically coupled power generators, but assume that the coupling is quasi-static and thus can be approximated as a constant in the diffeomorphism. In contrast, focusing on pure-feedback systems, we identify a class of dynamic couplings where the joint system is provably flat and explicitly construct its flat outputs and diffeomorphism, without assuming quasi-static coupling.
@@ III Flatness-Preserving Coupling
We denote the coupling term affecting subsystem $i$ as $\Delta^{i}$, and further, its effect on the $j^{\textrm{th}}$ substate as $\Delta^{i}_{j}$, such that
@@
A coupling term ${\Delta}(\mathbf{x})$ is lower-triangular on $\mathcal{X}$ if its effect on the $j^{\textrm{th}}$ substate of the subsystem $i$, $\Delta^{i}_{j}$, is smooth and can be expressed as
@@
While lower-order states from any subsystem may affect the $i$-th subsystem’s states of the same or higher order, the dependence on $\mathbf{x}^{<i}_{j+1}$ in (6) implicitly imposes an acyclic structure: higher-indexed subsystems may be affected by $(j+1)^{\textrm{th}}$ order states from lower-indexed ones, but not vice versa. The lower-triangular structure allows for a cascade decomposition of the joint system, which is closely related to the existence of invariant foliations on the state manifold (Fliess, 1985). For simplicity, we present Definition 1 and the subsequent results for the canonical case where this acyclic ordering is fixed across $\mathcal{X}$ and coincides with the subsystem indices. We leave extending our results to a case where the ordering is state-varying to future work.
We emphasize that the acyclicity restriction only applies to $(j+1)^{\textrm{th}}$ order dependencies. In Section IV, we analyze a special case, named strongly lower-triangular couplings (Def. 2), where such dependencies are absent . In this case, no acyclic ordering is required, allowing the framework to accommodate bidirectional and cyclic couplings. While the lower-triangular assumption does not cover all arbitrary coupling structures, it captures highly relevant physical interactions, e.g., networks of quadrotors dynamically coupled via downwash, where empirically, the dominant aerodynamic effects depend only on the relative position of the vehicles (Shi et al., 2021). In Figure 1, we show an illustration of these structural assumptions for a specific choice of $i$ and $j$.
@@
for all $(\mathbf{x},\mathbf{u})\in\mathcal{X}$} is known [^5] The existence of such maps are guaranteed by Assumption 1 and the implicit function theorem. (See Yang et al. (2025) for a proof.) For many well-known flat systems, they are known and can be expressed in closed-form..
@@
Under Assumption 2, the inverse maps $h_{k}^{i}$ are known in closed form. Consequently, for systems with moderate complexity, the maps $\Phi_{k}^{i}$ can often be explicitly derived offline, either by hand or via symbolic differentiation, thus incurring minimal computational overhead online. For systems with a high relative degree $r$ or highly complex coupling terms, the required time derivatives of $\Delta_{k-1}^{i}$ can be evaluated online by employing Taylor-mode automatic differentiation (AD), which allows the computational complexity to scale polynomially in $r$ while maintaining numerical stability.
@@ V Distributed Tracking Controller
where $K^{i}\in\mathbb{R}^{m_{i}\times rm_{i}}$ is a stabilizing feedback gain, which can be chosen, e.g., via pole placement. We can then compute the control input $\mathbf{u}^{i}$ by evaluating the flatness diffeomorphism $\Psi^{i}$, which depends on flat state $\mathbf{z}^{i}$, virtual input $\mathbf{v}^{i}$, and $(\mathbf{z}^{j},\mathbf{v}^{j})$ of an appropriate subset of the other subsystems. By the results in Section IV, subsystem $i$ needs to gather $(\mathbf{z}^{j},\mathbf{v}^{j})$ from subsystems $j\in\mathbf{Anc}_{\mathcal{G}(\mathbf{x})}(i)(\mathbf{x})$ to evaluate $\Psi^{i}$. Further, when the coupling is strongly lower-triangular, this dependency is further localized to its $r$-hop neighbors, $\mathcal{N}^{r}_{\mathcal{G}(\mathbf{x})}(i)$. In the prior case, discovering and routing data through a dynamic, global ancestor graph online can be difficult; thus, practical implementation might require the communication graph to be known and established a priori. However, in the latter case, since the coupling structure often reflects spatial proximity, agents can dynamically discover these dependencies online using localized peer-to-peer communication. Further, this allows for efficient distributed implementation as the information sharing between the systems does not scale with the number of subsystems $N$. The overall pipeline is summarized in Algorithm 1.
@@ VII Conclusion
We proposed a distributed controller for networks of flat, pure-feedback systems under dynamic coupling. We identified a class of lower-triangular couplings that preserve the flatness for the joint system and allow the joint flatness diffeomorphism to be explicitly constructed. We analyzed and exploited the sparsity pattern of the flatness diffeomorphism to design distributed controllers under communication constraints. We validated the framework in simulation for planar quadrotors coupled by downwash drag, showing that our flatness-based distributed controller achieves accurate tracking. Future work includes hardware experiments as well, relaxing the lower-triangular constraints on flatness-preserving couplings, and incorporating adaptive control techniques to reject disturbances that arise from inaccuracies in approximate coupling models (Join et al., 2024).